OpenStudy (pokemon23):

how do I solved this 4n= 4/n^2

OpenStudy (anonymous):

get n^2 out of the denominator by multiplying by n^2 4n^3 = 4 n^3 = 1 n = 1

OpenStudy (cathyangs):

you multiple both sides by n^2, making it 4n^3=4 the only possible answer is 1.

OpenStudy (jamesj):

yes exactly

OpenStudy (pokemon23):

hi

OpenStudy (cathyangs):

yo :D

OpenStudy (pokemon23):

I don't get it?

OpenStudy (pokemon23):